OpenStudy (anonymous):

help plz known function f: R..> R defined by \[f(x)= ax ^{3} + bx ^{2} + cx+d\] is that f (0) = 4 and its graph has an inflection point at (1,2). Knowing well that the tangent line to the graph of f at the point of abscissa x = 0 in horizontal. Calula a, b​​, c, and d

hartnn (hartnn):

ok, f(0)=4 from the equation, f(0)=... ? put x=0

OpenStudy (anonymous):

\[f(0) =4 ..> a*0^{3} + b*0^{2} c*0 +d =4 ...> d=4\] \[f''(1)=2 ..> 6a*1+2b=2..> 6a+2b=2.\] \[f'(0)=0 ..> 3*a*0^{2} +2b*0+c=0 ..> c=0\] is is correct?

hartnn (hartnn):

yes.

OpenStudy (anonymous):

how to get b

OpenStudy (anonymous):

b=0 and a =1/3 ???

OpenStudy (anonymous):

At an inflection point, the second derivative is 0. So f(1)=2, but f''(1)=0. You should have 6a=-2b, not 6a=-2b+2. Then at (1,2) you can say\[2=a \times 1^{3}+b \times 1^{2}+c \times x+d=a+b+c+d=a+b\]\[2=a+b=a-3a\]\[a=-1\]\[b=3\]

OpenStudy (anonymous):

c and d??

OpenStudy (anonymous):

@Aylin

OpenStudy (anonymous):

Didn't you already say that c=d=0?

OpenStudy (anonymous):

d=4 and c=0

OpenStudy (anonymous):

Ah, right. Then 2=a+b+c+d 2=a+b+d 2=a-3a+4 -2=-2a a=1 b=-3 c=0 d=4

OpenStudy (anonymous):

\[f''(1)=2..> 6ax+2b=2...> 6a+2b=2 ...> 6a+2*0=2 ..> a=1/3\] \[f''(0)=0 ..> 6a*0+2b=0 ..> 2b=0..> b=0\] i get a=1/3 and b =0

hartnn (hartnn):

f"(0)=0 is not correct!

OpenStudy (anonymous):

then

OpenStudy (anonymous):

The second derivative at (1,2) is equal to 0. The second derivative at (0,0) is not equal to 0.

OpenStudy (anonymous):

then f(0)= ??

OpenStudy (anonymous):

f(0)=4

OpenStudy (anonymous):

how you get b=3 i dont understand

OpenStudy (anonymous):

f''(1) should equal 0 since its an inflection point, not 2. f(1) = 2.

OpenStudy (anonymous):

\[f''(1)=2 ..> 6a+2b=2\] then ??

OpenStudy (anonymous):

No! \[f''(1)=0\]

OpenStudy (anonymous):

Since x = 1 is an inflection point.

OpenStudy (anonymous):

f''(1) =0 f''(2)=0

OpenStudy (anonymous):

\[f''(1)=0 ..> 6a+2b=0\] \[f''(2)=0 ..> 12a+2b=0\]

OpenStudy (anonymous):

You use f(0)=4 to find d. d=4 You use f'(0)=0 to find c. c=0 You use f''(1)=0 to find b in terms of a. You get 0=6a+2b -> b=-3a You use f(1)=2 to find out what a is. You get\[a+b+c+d=2\]\[c=0,d=4,b=-3a\]\[a-3a+4=2\]\[-2a=-2\]\[a=1\]Then you use b=-3a. a is 1, so b=-3 That is the entire problem.

OpenStudy (anonymous):

I don't know why you're trying to use f''(2)=0, since that isn't true.

OpenStudy (anonymous):

because (1,2) is inflection point

OpenStudy (anonymous):

Right, but that doesn't have ANYTHING to do with f''(2). What that means is that f''(1)=0

OpenStudy (anonymous):

ok